# Online Transfer Learning for RSV Case Detection

### 03 Feb 2024 v1

@@ VII Conclusion
In our study, we introduced an innovative online multi-source transfer learning methodology with a dynamic weighting mechanism and applied it to analyzing RSV data, systematically gathered over several seasons from actual patient encounters at UPMC. Our experimental results robustly demonstrate that this method significantly outperforms established baseline techniques, highlighting the considerable potential of our framework. Notably, it adeptly addresses the challenges inherent in biomedical research, such as the scarcity of labeled data and the complexities associated with sequential data. The success of our approach not only affirms its effectiveness but also paves the way for advanced predictive models in healthcare, offering a promising avenue for future explorations in the domain of sequential epidemiological data analysis.
@@ VIII Acknowledgements
This work was supported by the research grants R00LM013383 and R01LM013509 from the National Library of Medicine, National Institutes of Health. The research data were retrieved by the University of Pittsburgh’s Health Record Research Request (R3), which was supported by the National Institutes of Health through the grant UL1 TR001857. Dr. Harry Hochheiser was supported by MIDAS grant U24GM132013. Dr. Jessi Espino was partially supported by the National Center for Immunization and Respiratory Diseases grant 5U01IP001184. We are truly grateful for the invaluable assistance and guidance from Dr. Richard K. Zimmerman and Dr. Mary Patricia Nowalk, epidemiologists at the University of Pittsburgh; Dr. Milos Hauskrecht, a professor of computer science from the University of Pittsburgh.

### 07 Apr 2024 v2

@@ VII Conclusion
This paper introduces an online multi-source transfer learning methodology featuring a dynamic weighting mechanism, applied to the development of RSV case detection models using multiple years’ EHR data from the UPMC. Experimental results demonstrate that this method significantly outperforms many baseline techniques, highlighting the considerable potential of the approach. Notably, it adeptly addresses the challenges inherent in biomedical research, such as the scarcity of labeled data and the complexities associated with sequential data. The results of our investigation provide support for the effectiveness of our MSAW method, but also pave the way for studying more advanced classification models for sequential data analysis.
@@ VIII Acknowledgements
This work was supported by the research grants R00LM013383 and R01LM013509 from the National Library of Medicine, National Institutes of Health (NIH). The research data were retrieved by the University of Pittsburgh’s Health Record Research Request, which was supported by the NIH grant UL1TR001857. Dr. Harry Hochheiser was supported by MIDAS grant U24GM132013. Dr. Jessi Espino was partially supported by the National Center for Immunization and Respiratory Diseases grant 5U01IP001184. We are grateful for the invaluable assistance and guidance from Dr. Richard K. Zimmerman and Dr. Mary Patricia Nowalk, and Dr. Milos Hauskrecht, from the University of Pittsburgh.